OpenStudy (anonymous):

If u know this u r a genius A coin is selected at random from a bag of money. Use the distribution table above to determine the expected value. Round the answer to the nearest cent. Coin COUNT Penny 17 Nickel 11 Dime 8 Quarter 14

OpenStudy (anonymous):

A. $0.05 B. $0.07 C. $0.10 D. $0.12

OpenStudy (anonymous):

there are a couple ways to do this perhaps the easiest was is to add up all the money, then divide by the number of coins there are

OpenStudy (anonymous):

in other words, it is really just the "average"

OpenStudy (anonymous):

so 12?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

just like the last one lets count the money first

OpenStudy (misty1212):

\[17+5\times 11+8\times 10+14\times 25\]

OpenStudy (misty1212):

i get \(502\)

OpenStudy (misty1212):

then divide that by the number of coins

OpenStudy (misty1212):

i get 50

OpenStudy (misty1212):

final answer \[\frac{502}{50}\] in pennies